OpenStudy (anonymous):

Slope of (5/2,1),(3,1/4)

OpenStudy (anonymous):

\[\frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (anonymous):

I get 3/4 / -2 and that can't be right

OpenStudy (anonymous):

\[\frac{ \frac{ 1}{ 4 } - 1}{3 - \frac{ 5 }{ 2 } }\]

OpenStudy (amistre64):

you can scale the points to eliminate the fractions ... id scale them by 4 4* (5/2,1) = (10,4) 4*(3,1/4) = (12,1)